OpenStudy (anonymous):

Solve: log base 2 (s^2 - 16) = 1 + log base 2 (s - 4)

OpenStudy (lgbasallote):

log base 2 (s^2 - 16) - log base 2 (s - 4) = 1

OpenStudy (kropot72):

Put all the log terms on one side. Can you do that as a first step?

OpenStudy (lgbasallote):

log base 2 (s^2 - 16)/(s-4) = 1 log base 2 (s+4)(s-4)/(s-4) = 1 log base 2 (s-4) = 1 change to exponential 2^1 = s - 4 2 = s - 4 2 + 4 = s 6 = s

OpenStudy (anonymous):

Hm.. my teacher gave us the answers so we could solve them out, and the answer to this one is no solution? o:

OpenStudy (lgbasallote):

oops made a mistake

OpenStudy (lgbasallote):

log base 2 (s+4) = 1 not s - 4 sorry 2^1 = s+4 2 - 4 = s -2 = s how is that no solution? o.O

OpenStudy (anonymous):

Oh because you can't have negatives O: pretty sure

OpenStudy (lgbasallote):

maybe...but it doesnt make sense...im thinking your teacher couldve made a mistake :/

OpenStudy (anonymous):

All I remember is something about logs and not being allowed to have negatives.. hmm. o;

OpenStudy (lgbasallote):

\[\large \log_2 (s^2 - 16) = 1 + \log_2 (s-4)\] this is the question right??

OpenStudy (lgbasallote):

yeah logs cant have negatives

OpenStudy (lgbasallote):

oh...oh yeah

OpenStudy (anonymous):

Yeahyeah, lol. c:

OpenStudy (lgbasallote):

if you sub -2 into \(\log_2 (s^2 - 16)\) you'll get a negative number so it's no solution

OpenStudy (anonymous):

Ahh, okay. Thank youu C: